What amount of thermal energy (in kJ) is required to convert 27.0 g of ethanol at -158 °C completely to gaseous ethanol at 85 °C? The melting point of ethanol is -114 °C and its normal boiling point is 78 °C.

The heat of fusion of ethanol is 5.0 kJ mol-1

The heat of vaporization of ethanol at its normal boiling point is 35.0 kJ mol-1

The specific heat capacity of solid ethanol is 1.05 J g-1 °C-1

The specific heat capacity of liquid ethanol is 2.44 J g-1 °C-1

The specific heat capacity of gaseous ethanol is 1.01 J g-1 °C-1
1
Expert's answer
2019-11-27T07:12:33-0500

E=(27*(1.05*(-114-(-158))+2.44*(78-(-114))+1.01*(85-78))+(27/46)*(35+5))/1000=14.110 (kJ)

Answer: 14.110 kJ.